# Introspectively Envy-Free and Efficient Allocation of Indivisible Mixed Manna

### v2

@@ 6.4 Proof of Theorem 13
Output : Allocation $\widehat{\mathcal{S}}=(\widehat{S}_{1},\dots,\widehat{S}_{n})\in\mathcal{M}^{*}$ satisfying $\varphi^{+}(\widehat{\mathcal{S}})>\varphi^{+}(\mathcal{S})$
@@ 8 $\mathrm{EF}$ and $\mathrm{fPO}$ Allocations for Divisible Mixed Manna
The above-mentioned characterization of $\mathrm{fPO}$ implies that the set all $\mathrm{fPO}$ fractional allocations $\mathscr{F}=\bigcup\limits_{\lambda\in\mathbb{R}^{n}_{>0}}\ \mathcal{F}_{\lambda}$. Note that this union is composed of finitely many faces; the polytope $\mathcal{F}$ itself has finite number of faces. Since every face of $\mathcal{F}$ is closed, we get that $\mathscr{F}$ is a finite union of closed sets, and hence, is closed as well. The lemma stands proved.
@@ Appendix A $\mathrm{EF}1$ and $\mathrm{PO}$ for Indivisible Goods
In this section, we sketch how the proof of Theorem 1 can be modified to establish the existence of $\mathrm{PO}$ and $\mathrm{EF}1$ allocations of indivisible goods. In particular, we will show the existence of allocations that are $\mathrm{PO}$, and extrospectively envy free up to one item $\mathrm{XEF}1$, which is defined below.
@@
An allocation $\mathcal{A}=(A_{1},A_{2},\ldots,A_{n})$ is called extrospectively envy free up to one item ($\mathrm{XEF}1$) for an instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$ if for all agents $i,j\in[n]$, there exists a subset $S\subseteq[m]$ with $|S|\leq 1$ such that $v_{i}(A_{i})\geq v_{i}(A_{j}\triangle S)$.
Note that for instances with only goods, $\mathrm{XEF}1$ coincides with $\mathrm{EF}1$. To avoid duplication, we will only focus on the changes required in the proof of Theorem 1 to establish the existence of $\mathrm{PO}$ and $\mathrm{XEF}1$ allocations.
The first change is in the definition of the primal-dual linear programs. We will denote these modified linear programs as $\overline{\textbf{{LP}}}$ and $\overline{\textbf{{Dual-LP}}}$.
@@
Here the definition of the non-degenerate instance, $\widehat{\mathcal{I}}$, and the valuations $\widehat{v}_{i}(j)$ are unchanged, follow Section 4.1. The following lemmas, which correspond to Lemmas 4 and 5, follow from an almost identical argument.
@@
If an allocation $\mathcal{A}=(A_{1},A_{2},\ldots,A_{n})$ is $\mathrm{XEF}1$ in the non-degenerate instance $\overline{\mathcal{I}}=\langle[n],[m+1],\{\widehat{v}_{i}\}_{i=1}^{n}\rangle$, then the allocation $(A_{1}\setminus\{m+1\},A_{2}\setminus\{m+1\},\ldots,A_{n}\setminus\{m+1\})$ is $\mathrm{XEF}1$ in the instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$.
@@
If an allocation $\mathcal{A}=(A_{1},A_{2},\ldots,A_{n})\in\mathrm{OPT}(\overline{\textbf{{LP}}}(w,\eta))$, for any $w\in\Delta_{n-1}$, then the allocation $(A_{1}\setminus\{m+1\},A_{2}\setminus\{m+1\},\ldots,A_{n}\setminus\{m+1\})$ is $\mathrm{PO}$ for the instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$.
@@
The second key change is how we apply the KKM theorem. We define the sets $\overline{C}_{1},\overline{C}_{2},\ldots,\overline{C}_{n}\subseteq\Delta_{n-1}$ by using $\operatorname*{arg\,min}$ instead of $\operatorname*{arg\,max}$ (as done in Equation 31) as shown below.
@@
Here $\overline{C}_{i}$ corresponds to the set of weights where agent $i$ can be made the worst, with respect to the bundle prices. This is counterintuitive and notably different from Section 5. The following lemma follows from an argument similar to Lemma 8.
@@
The set $\overline{C}_{i}\subseteq\Delta_{n-1}$ is closed, for each $i\in[n]$.
We will now state and prove the following counterpart of Lemma 9; its proof has notable differences as compared to Lemma 9.
@@
For any $w\in\Delta_{n-1}$ there exists an $i\in\mathrm{supp}(w)$, such that $w\in\overline{C}_{i}$.
@@
Case I: $\mathrm{supp}(w)=[n]$. Consider any allocation $\mathcal{A}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(w,\eta))$ and $p=\mathrm{OPT}(\overline{\textbf{{Dual-LP}}}(w,\eta))$. Let $i^{\prime}\in\operatorname*{arg\,min}_{j\in[n]}p(A_{j})$. Since $\mathrm{supp}(w)=[n]$, we have $i^{\prime}\in\mathrm{supp}(w)$. Additionally, we have $w\in\overline{C}_{i^{\prime}}$ since $i^{\prime}\in\operatorname*{arg\,min}_{j\in[n]}p(A_{j})$ where $\mathcal{A}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(w,\eta))$.
Case II: $\mathrm{supp}(w)\subsetneq[n]$. In this case we have $n]\setminus\mathrm{supp}(w)\neq\emptyset$. Define $\{i\in[n]\mid w_{i}\geq 1/n\}; HH\neq\emptyset$ since $\sum_{i=1}^{n}w_{i}=1$. The set $[n]$ can be partitioned into three disjoint sets $n]\setminus\mathrm{supp}(w)$, $H$, and $\mathrm{supp}(w)\setminus H$. Let prices $p=\mathrm{OPT}(\overline{\textbf{{Dual-LP}}}(w,\eta))$ and $\mathcal{A}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(w,\eta))$ be any allocation. Let $h\in\operatorname*{arg\,max}_{i\in H}w_{i}$ be an agent having the highest corresponding weight value; note that $h\in H\subseteq\mathrm{supp}(w)$.
To show that there exists an $i^{\prime}\in\mathrm{supp}(w)$ such that $w\in\overline{C}_{i^{\prime}}$, we will first show that $\max\limits_{i\in[n]\setminus\mathrm{supp}(w)}p(A_{i})\geq p(A_{\ell})$. This will conclude the proof, as we will get
$\max\limits_{i\in[n]\setminus\mathrm{supp}(w)}p(A_{i})\geq p(A_{\ell})\geq\min\limits_{i\in\mathrm{supp}(w)}p(A_{i}),$
where the final inequality follows from $\ell\in\mathrm{supp}(w)$. This will in turn imply the existence of an $i^{\prime}\in\mathrm{supp}(w)$ such that $i^{\prime}\in\operatorname*{arg\,min}_{j\in[n]}p(A_{j})$, i.e., $w\in\overline{C}_{i^{\prime}}$, completing the proof.
Hence, in the subsequent proof, we will show that $\max\limits_{i\in[n]\setminus\mathrm{supp}(w)}p(A_{i})\geq p(A_{\ell})$. Towards this, we will prove the following inequalities,
$\max\limits_{i\in[n]\setminus\mathrm{supp}(w)}p(A_{i})\geq\lambda/2\eta$ (32)
$\lambda/2\eta\geq p(A_{\ell})$ (33)
Together $(i)$ and $(ii)$ will imply the desired inequality, $\max\limits_{i\in[n]\setminus\mathrm{supp}(w)}p(A_{i})\geq\lambda/2\eta\geq p(A_{\ell})$.
@@
We will now show that no chore $c\in C$ is allocated to the agent $i\in[n]\setminus\mathrm{supp}(w)$. To this end, it is sufficient to prove $\widehat{v}_{\ell}(c)/(w_{\ell}+\eta)>\widehat{v}_{i}(c)/\eta$ or equivalently $\eta\widehat{v}_{\ell}(c)>(w_{\ell}+\eta)\widehat{v}_{i}(c)$. To establish this, we consider,
$(w_{\ell}+\eta)|\widehat{v}_{i}(c)|$ $>|\widehat{v}_{i}(c)|/n$ ($w_{\ell}\geq 1/n$, $\eta>0$)
@@
$>\eta\max_{a,b}|\widehat{v}_{a}(b)|$ (Equation 4)
$>\eta|\widehat{v}_{\ell}(c)|.$
Since $\widehat{v}_{\ell}(c),\widehat{v}_{i}(c)<0$, we get the desired inequality $\eta\widehat{v}_{\ell}(c)>(w_{\ell}+\eta)\widehat{v}_{i}(c)$.
Establishing Inequality (33). If an item $t\in[m]$ is allocated to agent $\ell$, then we have $p_{t}=\widehat{v}_{\ell}(t)/(w_{\ell}+\eta)\leq\widehat{v}_{\ell}(t)/w_{\ell}\leq n\widehat{v}_{\ell}(t)\leq n\max_{a,b}|{\widehat{v}}_{a}(b)|$, where the second inequality follows from $w_{\ell}\geq 1/n$. Therefore, we get, $p(A_{\ell})\leq\sum_{t\in A_{\ell}}p_{t}\leq m\cdot n\max_{a,b}|{\widehat{v}}_{a}(b)|$. Finally, using Equation 4 gives us Inequality (33), $p(A_{\ell})\leq m\cdot n\max_{a,b}|{\widehat{v}}_{a}(b)|=\lambda/2\eta$.

### v3

@@ 6.4 Proof of Theorem 13
Output : Allocation $\widehat{\mathcal{S}}=(\widehat{S}_{1},\dots,\widehat{S}_{n})\in\mathcal{M}^{*}$ satisfying $\varphi^{+}(\widehat{\mathcal{S}})>\varphi^{+}(\mathcal{S})$.
@@ 8 $\mathrm{EF}$ and $\mathrm{fPO}$ Allocations for Divisible Mixed Manna
The above-mentioned characterization of $\mathrm{fPO}$ implies that the set all $\mathrm{fPO}$ fractional allocations $\mathscr{F}=\bigcup\limits_{\lambda\in\mathbb{R}^{n}_{>0}}\ \mathcal{F}_{\lambda}$. Note that this union is composed of finitely many faces; the polytope $\mathcal{F}$ itself has finite number of faces. Since every face of $\mathcal{F}$ is closed, we get that $\mathscr{F}$ is a finite union of closed sets, and hence, is closed as well. The lemma stands proved. ∎
@@ Appendix A $\mathrm{EF}1$ and $\mathrm{PO}$ for Indivisible Goods
This section provides a sketch of how the proof of Theorem 1 can be modified to establish the existence of $\mathrm{PO}$ and $\mathrm{EF}1$ allocations of indivisible goods. In particular, we will show the existence of mixed-manna allocations that are $\mathrm{PO}$, and extrospectively envy-free up to one item ($\mathrm{XEF}1$); this fairness notion is defined next.
@@
In an instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$, an allocation $\mathcal{A}=(A_{1},\ldots,A_{n})$ is said to be extrospectively envy-free up to one item ($\mathrm{XEF}1$) if, for all agents $i,j\in[n]$, there exists a subset $S\subseteq[m]$ of size at most one ($|S|\leq 1$) such that $v_{i}(A_{i})\geq v_{i}(A_{j}\triangle S)$.
To avoid repeating the arguments used in the proof of Theorem 1, we will primarily focus on the modifications required to establish the existence of $\mathrm{PO}$ and $\mathrm{XEF}1$ allocations.
To prove this existential result, some changes are required in the construction of the non-degenerate instance $\widehat{\mathcal{I}}=\langle[n],[\widehat{m}],\{\widehat{v}_{i}\}_{i=1}^{m}\rangle$ (see Section 4.1 for comparison) and in the primal and dual programs (see Section 4.2). We begin by describing the changes in $\widehat{\mathcal{I}}$; with a slight abuse of notations we continue to denote the non-degenerate instance by $\widehat{\mathcal{I}}$. We set $\widehat{m}\coloneqq m+n$, i.e., we add $n$ auxiliary items to the given instance $\mathcal{I}$. For each $i\in[n]$, the valuations of the new item, $(m+i)$, is set as follows: $\widehat{v}_{i}(m+i)=\lambda/2$ and $\widehat{v}_{j}(m+i)=0$, for all $j\in[n]\setminus\{i\}$. For the given items $t\in[m]$ and any agent $i\in[n]$, if $v_{i}(t)=0$, we set $\widehat{v}_{i}(t)=0$, i.e., leave the zero values unchanged. Otherwise, if $v_{i}(t)\neq 0$, then we set $\widehat{v}_{i}(t)=v_{i}(t)-\overline{\varepsilon}_{i,t}$. The perturbations $\overline{\varepsilon}_{i,t}$ are drawn independently from the uniform distribution $\mathrm{Uniform}[0,\overline{\varepsilon}]$, where $\overline{\varepsilon}>0$ is fixed to satisfy
$\overline{\varepsilon}<\frac{\lambda\omega}{16mn^{2}(m+n)\ \max\limits_{i\in[n],t\in[m]}|v_{i}(t)|}$ (31)
Here, $\lambda$ and $\omega$ are defined as in Section 4.1. We will now describe the modified linear programs, denoted by $\overline{\textbf{{LP}}}$ and $\overline{\textbf{{Dual-LP}}}$. Their construction will utilize the following constant
$\overline{\eta}\coloneqq\frac{\lambda}{2n(m+n)\max\limits_{i\in[n],\ t\in[\widehat{m}]}|\widehat{v}_{i}(t)|}$ (32)
The constants $\overline{\varepsilon}$ and $\overline{\eta}$ satisfy the following analogue of Equation 5
$\overline{\varepsilon}<\frac{\overline{\eta}\omega}{4mn}$ (33)
@@
The following lemma is a counterpart of Lemma 4; its proof follows similar arguments and is omitted.
@@
If an allocation $\mathcal{A}=(A_{1},\ldots,A_{n})$ is $\mathrm{XEF}1$ in the non-degenerate instance $\widehat{\mathcal{I}}=\langle[n],[\widehat{m}],\{\widehat{v}_{i}\}_{i=1}^{n}\rangle$, then the allocation $(A_{1}\setminus\{(m+i)\}_{i=1}^{n},A_{2}\setminus\{(m+i)\}_{i=1}^{n},\ldots,A_{n}\setminus\{(m+i)\}_{i=1}^{n})$ is $\mathrm{XEF}1$ in the given instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$.
We will now state an analogue of Lemma 5. While the proof here follows an approach similar to that of Lemma 5, key technical aspects differ and, hence, we include the proof below.
@@
If an allocation $\mathcal{A}=(A_{1},\ldots,A_{n})\in\mathrm{OPT}(\overline{\textbf{{LP}}}(w,\overline{\eta}))$, for any $w\in\Delta_{n-1}$, then the allocation $(A_{1}\setminus\{(m+i)\}_{i=1}^{n},A_{2}\setminus\{(m+i)\}_{i=1}^{n},\ldots,A_{n}\setminus\{(m+i)\}_{i=1}^{n})$ is $\mathrm{PO}$ in the given instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$.
Proof.
Define an intermediate instance $\mathcal{I}^{\prime}=\langle[n],[m+n],\{v^{\prime}_{i}\}_{i=1}^{n}\rangle$ wherein $v^{\prime}_{i}(t)=v_{i}(t)$, for all agents $i\in[n]$ and given items $t\in[m]$. In addition, set $v^{\prime}_{i}(m+t)=\widehat{v}_{i}(m+t)$ for each $i,t\in[n]$. Note that instance $\mathcal{I}^{\prime}$ is constructed by adding the auxiliary items $m+1,\ldots,m+n$ to $\mathcal{I}$ and keeping the valuations of the original $m$ items unchanged. Instance $\mathcal{I}^{\prime}$ is defined for the purposes of analysis: We will first show that $\mathcal{A}$ is $\mathrm{PO}$ in $\mathcal{I}^{\prime}$. Then, we will conclude the proof by showing that $(A_{1}\setminus\{(m+i)\}_{i=1}^{n},\ldots,A_{n}\setminus\{(m+i)\}_{i=1}^{n})$ is $\mathrm{PO}$ in the given instance $\mathcal{I}$.
Towards a contradiction, we assume that there exists an allocation $\mathcal{B}=(B_{1},\ldots,B_{n})$ that Pareto dominates $\mathcal{A}$ in $\mathcal{I}^{\prime}$. That is, $v^{\prime}_{i}(B_{i})\geq v^{\prime}_{i}(A_{i})$, for all $i\in[n]$, and $v^{\prime}_{a}(B_{a})>v^{\prime}_{a}(A_{a})$ for some $a\in[n]$. Without loss of generality, we also assume that $\mathcal{B}$ is Pareto efficient in $\mathcal{I}^{\prime}$; this ensures that, for each $i\in[n]$, the auxiliary item $(m+i)$ is allocated to agent $i$ in $\mathcal{B}$. Additionally, note that each auxiliary item $(m+i)$ also gets allocated to agent $i$ in $\mathcal{A}$ since $\mathcal{A}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(w,\overline{\eta}))$. These observations imply that
$0$ $<\sum_{i=1}^{n}v^{\prime}_{i}(B_{i})-\sum_{i=1}^{n}v^{\prime}_{i}(A_{i})$
$=n\frac{\lambda}{2}+\sum_{i=1}^{n}v^{\prime}_{i}(B_{i}\setminus\{(m+i)\}_{i=1}^{n})-n\frac{\lambda}{2}-\sum_{i=1}^{n}v^{\prime}_{i}(A_{i}\setminus\{(m+i)\}_{i=1}^{n})$ ($v^{\prime}_{i}(m+i)=\lambda/2$ for all $i$)
$=\sum_{i=1}^{n}v_{i}(B_{i}\setminus\{(m+i)\}_{i=1}^{n})-\sum_{i=1}^{n}v_{i}(A_{i}\setminus\{(m+i)\}_{i=1}^{n})$ (34)
That is, $\sum_{i=1}^{n}v_{i}(B_{i}\setminus\{(m+i)\}_{i=1}^{n})-\sum_{i=1}^{n}v_{i}(A_{i}\setminus\{(m+i)\}_{i=1}^{n})>0$. Using this strict inequality and the definition of $\omega$, we obtain
$\omega\leq\sum_{i=1}^{n}v_{i}(B_{i}\setminus\{m+i\}_{i=1}^{n})-\sum_{i=1}^{n}v_{i}(A_{i}\setminus\{m+i\}_{i=1}^{n})=\sum_{i=1}^{n}v^{\prime}_{i}(B_{i})-\sum_{i=1}^{n}v^{\prime}_{i}(A_{i})$
Next, we use the welfare optimality of $\mathcal{A}$ in $\widehat{\mathcal{I}}$ (and under the weight vector $w$) to arrive at a contradiction. In particular, $\mathcal{A}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(w,\overline{\eta}))$ implies that $\sum_{i=1}^{n}(w_{i}+\overline{\eta})^{-1}\widehat{v}_{i}(B_{i})\leq\sum_{i=1}^{n}(w_{i}+\overline{\eta})^{-1}\widehat{v}_{i}(A_{i})$. For each subset $S\subseteq[\widehat{m}$] and each agent $i$, it holds that $v^{\prime}_{i}(S)-\overline{\varepsilon}|S|\leq\widehat{v}_{i}(S)\leq v^{\prime}_{i}(S)$. Therefore, we have $\sum_{i=1}^{n}(w_{i}+\overline{\eta})^{-1}(v^{\prime}_{i}(B_{i})-\overline{\varepsilon}|B_{i}|)\leq\sum_{i=1}^{n}(w_{i}+\overline{\eta})^{-1}v^{\prime}_{i}(A_{i})$. This inequality reduces to
$\sum_{i=1}^{n}\frac{1}{(w_{i}+\overline{\eta})}(v^{\prime}_{i}(B_{i})-v^{\prime}_{i}(A_{i}))$ $\leq\overline{\varepsilon}\sum_{i=1}^{n}\frac{1}{(w_{i}+\overline{\eta})}|B_{i}|$
$\leq\overline{\varepsilon}\sum_{i=1}^{n}\frac{1}{\overline{\eta}}|B_{i}|$ ($w_{i}\geq 0$)
$=\frac{\overline{\varepsilon}(m+n)}{\overline{\eta}}$ (36)
We can lower bound the left-hand-side of inequality (36) as follows
$\sum_{i=1}^{n}\frac{1}{(w_{i}+\overline{\eta})}(v^{\prime}_{i}(B_{i})-v^{\prime}_{i}(A_{i}))$ $\geq\frac{1}{2}\sum_{i=1}^{n}(v^{\prime}_{i}(B_{i})-v^{\prime}_{i}(A_{i}))$ ($v^{\prime}_{i}(B_{i})-v^{\prime}_{i}(A_{i})\geq 0$ and $w_{i}+\overline{\eta}\leq 2$)
$\geq\frac{\omega}{2}$ (via (A))
The last inequality and equation (36) imply $\frac{\omega}{2}\leq\frac{\overline{\varepsilon}(m+n)}{\overline{\eta}}$, or equivalently, $\frac{\overline{\eta}\omega}{2(m+n)}\leq\overline{\varepsilon}$. This bound, however, contradicts the strict inequality (33), which states that $\overline{\varepsilon}<\frac{\overline{\eta}\omega}{4mn}$. [^15] We have $n,m\geq 2$. Therefore, by way of contradiction, we obtain that no allocation $\mathcal{B}$ Pareto dominates $\mathcal{A}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(w,\overline{\eta}))$ in the instance $\mathcal{I}^{\prime}$, i.e., $\mathcal{A}$ is $\mathrm{PO}$ in $\mathcal{I}^{\prime}$.
Finally, we will use the Pareto optimality of $\mathcal{A}$ in $\mathcal{I}^{\prime}$ to show that the allocation $(A_{1}\setminus\{(m+i)\}_{i=1}^{n},\ldots,A_{n}\setminus\{(m+i)\}_{i=1}^{n})$ is $\mathrm{PO}$ in the given instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$. Assume, towards a contradiction, that allocation $(D_{1},\ldots,D_{n})$ Pareto dominates $(A_{1}\setminus\{(m+i)\}_{i=1}^{n},\ldots,A_{n}\setminus\{(m+i)\}_{i=1}^{n})$ in $\mathcal{I}$. As observed above, for each $i\in[n]$, the item $(m+i)$ is allocated to agent $i$ in allocation $\mathcal{A}$. Hence, our assumption would imply that $(D_{1}\cup\{(m+1)\},\ldots,D_{i}\cup\{(m+i)\},\ldots,D_{n}\cup\{(m+n)\})$ Pareto dominates $(A_{1},\ldots,A_{i},\ldots,A_{n})$ in $\mathcal{I}^{\prime}$. Since this contradicts the already established Pareto optimality of $\mathcal{A}$ in $\mathcal{I}^{\prime}$, we obtain that $(A_{1}\setminus\{(m+i)\}_{i=1}^{n},\ldots,A_{n}\setminus\{(m+i)\}_{i=1}^{n})$ is $\mathrm{PO}$ in the instance $\mathcal{I}$. The lemma stands proved. ∎
@@
Another key change is how we apply the KKM theorem. Here, we define the sets $\overline{C}_{1},\overline{C}_{2},\ldots,\overline{C}_{n}\subseteq\Delta_{n-1}$ considering $\operatorname*{arg\,min}$ over the optimal prices (instead of $\operatorname*{arg\,max}$ as in Equation 10).
@@
Note that each $\overline{C}_{i}$ is the set of weights $w\in\Delta_{n-1}$ for which there is an optimal allocation $\mathcal{A}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(w,\overline{\eta}))$ wherein $i$’s bundle has the lowest price, under the optimal $p$. This defining condition, that agent $i$ is worst off in terms of optimal bundle prices, is in complete contrast to the one imposed in Section 5 (Equation 10). Nonetheless, we show below that this somewhat counterintuitive condition leads to the desired $\mathrm{XEF}1$ guarantee.
The following lemma asserts that the sets $\overline{C}_{1},\ldots,\overline{C}_{n}\subseteq\Delta_{n-1}$ are closed; the proof of this lemma is similar to that of Lemma 8.
@@
For each $i\in[n]$, the set $\overline{C}_{i}\subseteq\Delta_{n-1}$ is closed.
We will now prove that the sets $\overline{C}_{1},\ldots,\overline{C}_{n}$ satisfy the KKM covering condition; the proof of this lemma is notably different from its counterpart, Lemma 9.
@@
For any $w\in\Delta_{n-1}$ there exists an $i\in\mathrm{supp}(w)$ such that $w\in\overline{C}_{i}$.
@@
Case I: $\mathrm{supp}(w)=[n]$. Consider any allocation $\mathcal{A}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(w,\overline{\eta}))$ and optimal prices $p=\mathrm{OPT}(\overline{\textbf{{Dual-LP}}}(w,\overline{\eta}))$. Let $i^{\prime}\in\operatorname*{arg\,min}_{j\in[n]}p(A_{j})$. In the current case $\mathrm{supp}(w)=[n]$ and, hence, $i^{\prime}\in\mathrm{supp}(w)$. Additionally, we have $w\in\overline{C}_{i^{\prime}}$, since $i^{\prime}\in\operatorname*{arg\,min}_{j\in[n]}p(A_{j})$. Hence, for vectors $w$ with full support, the KKM covering condition holds.
Case II: $\mathrm{supp}(w)\subsetneq[n]$. In this case $n]\setminus\mathrm{supp}(w)\neq\emptyset$. Write optimal prices $p=\mathrm{OPT}(\overline{\textbf{{Dual-LP}}}(w,\overline{\eta}))$ and consider any allocation $\mathcal{A}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(w,\overline{\eta}))$. We will show that there exists an $i^{\prime}\in\mathrm{supp}(w)$ with the property that $i^{\prime}\in\operatorname*{arg\,min}_{j\in[n]}\ p(A_{j})$. For such an agent $i^{\prime}$, we have $w\in C_{i^{\prime}}$, and, hence, we will obtain KKM covering in the current case as well.
Define $\coloneqq\{i\in[n]\mid w_{i}\geq 1/n\}; HH\neq\emptyset$ since $\sum_{i=1}^{n}w_{i}=1$. Also, the set of agents $[n]$ partitions into three disjoint subsets: $n]\setminus\mathrm{supp}(w)$, $H$, and $\mathrm{supp}(w)\setminus H$.
Let $h\in\operatorname*{arg\,max}_{i\in H}w_{i}$ be an agent with the highest weight value; note that $h\in H\subseteq\mathrm{supp}(w)$.
To show that there exists an $i^{\prime}\in\mathrm{supp}(w)$ such that $w\in\overline{C}_{i^{\prime}}$, we will first show that $\min\limits_{i\in[n]\setminus\mathrm{supp}(w)}p(A_{i})\geq p(A_{h})$. This will conclude the proof, since we will obtain
$\min\limits_{i\in[n]\setminus\mathrm{supp}(w)}p(A_{i})\geq p(A_{h})\geq\min\limits_{j\in\mathrm{supp}(w)}p(A_{j}).$
Here, the final inequality follows from $h\in\mathrm{supp}(w)$. This will in turn imply the existence of an $i^{\prime}\in\mathrm{supp}(w)$ such that $i^{\prime}\in\operatorname*{arg\,min}_{j\in[n]}p(A_{j})$, i.e., $w\in\overline{C}_{i^{\prime}}$, completing the proof.
Hence, in the remainder of the proof, we will show that $\min\limits_{i\in[n]\setminus\mathrm{supp}(w)}p(A_{i})\geq p(A_{h})$. Towards this, we will prove the following two inequalities
$\min\limits_{i\in[n]\setminus\mathrm{supp}(w)}p(A_{i})\geq\lambda/2\overline{\eta}$ (38)
$\lambda/2\overline{\eta}\geq p(A_{h})$ (39)
Together inequalities (38) and (39) imply the desired bound $\min\limits_{i\in[n]\setminus\mathrm{supp}(w)}p(A_{i})\geq\lambda/2\overline{\eta}\geq p(A_{h})$.
@@
We will now show that no chore $c\in C$ is allocated to the agent $i\in[n]\setminus\mathrm{supp}(w)$. To this end, it is suffices to show that $\widehat{v}_{h}(c)/(w_{h}+\overline{\eta})>\widehat{v}_{i}(c)/\overline{\eta}$, or equivalently $\overline{\eta}\widehat{v}_{h}(c)>(w_{h}+\overline{\eta})\widehat{v}_{i}(c)$. To establish this, note that
$(w_{h}+\overline{\eta})|\widehat{v}_{i}(c)|$ $>|\widehat{v}_{i}(c)|/n$ ($w_{h}\geq 1/n$ and $\overline{\eta}>0$)
@@
$>\overline{\eta}\ \max_{a,b}|\widehat{v}_{a}(b)|$ (Equation 32)
$\geq\overline{\eta}\ |\widehat{v}_{h}(c)|.$
Since $\widehat{v}_{h}(c),\widehat{v}_{i}(c)<0$, we get the desired inequality $\overline{\eta}\widehat{v}_{h}(c)>(w_{h}+\overline{\eta})\widehat{v}_{i}(c)$.
Establishing Inequality (39). If an item $t\in[\widehat{m}$] is allocated to agent $h$, then we have $p_{t}=\widehat{v}_{h}(t)/(w_{h}+\overline{\eta})\leq\max_{a,b}|{\widehat{v}}_{a}(b)|/(w_{h}+\overline{\eta})<\max_{a,b}|{\widehat{v}}_{a}(b)|/w_{h}\leq n\max_{a,b}|{\widehat{v}}_{a}(b)|$, where the final inequality follows from $w_{h}\geq 1/n$. Therefore, $p(A_{h})\leq\sum_{t\in A_{h}}p_{t}\leq(m+n)\ n\max_{a,b}|{\widehat{v}}_{a}(b)|$. Finally, using the definition of $\overline{\eta}$ (Equation 32), we obtain the inequality (39): $p(A_{h})\leq n(m+n)\max_{a,b}|{\widehat{v}}_{a}(b)|=\lambda/2\overline{\eta}$.
This completes the proof of the lemma. ∎